a minor league baseball team plays 130 games in a season. if the team won 14 more than three times as many games as they lost, how many wins and losses did the team have?
Answer by palanisamy(496)

Let the number of games won = x
Let the number of games lost = y
Given, x+y = 130 ...(1)
Next, x = 3y+14
x-3y = 14 ...(2)
(1)-(2)=> 4y=116
y = 116/4
y = 29
(1)=> x+29 = 130
x=130-29
x=101